# DMesh++: An Efficient Differentiable Mesh for Complex Shapes

### v1

@@ 4.4 Loss Functions
Figure 6: Comparison of tessellation speed. Our method based on minimum-ball algorithm computes face probabilities up to 16 times (2D) and 32 times (3D) faster than DMesh [37].
Figure 7: Quality of DMesh++ reconstructed from 2D point cloud sampled from Egyptian painting. In left part, “imaginary” part (gray) and “real” part (blue) are rendered together.
@@ 5 Experiments
This section presents our experimental results. First, we evaluate how Minimum-Ball accelerates the tessellation function in 2D and 3D. Next, we demonstrate the Reinforce-Ball algorithm’s effectiveness in producing efficient 2D meshes with adaptive resolution. Finally, we showcase a practical application in 3D multi-view reconstruction, illustrating that our method is suitable for (differentiable) downstream tasks involving complex shapes. Our main algorithms are implemented in PyTorch [31], with a differentiable 3D renderer in CUDA [26]. All experiments were conducted on a system with an AMD EPYC 7R32 CPU and Nvidia A10 GPU.
DMesh [37] (0) 1.97 2506 2245 30.39
DMesh ($10^{-4}$) 2.68 666 693 153.10
DMesh ($10^{-3}$) 12.48 456 488 152.37
DMesh++ (0) 1.82 2862 2793 11.33
DMesh++ ($10^{-6}$) 1.86 1386 1394 278.88
DMesh++ ($10^{-5}$) 2.77 149 152 200.05
5.1 Tessellation Speed
Figure 9: Qualitative comparison of 3D multi-view reconstruction results for open surface. Here we illustrate from back of an open surface model (the front view is rendered at the left top of (a)). Colors represent inside and outside facing surfaces. DMesh++ captures geometric details better than the other methods, while GShell [19] suffers from intricate structure of the shape, and DMesh [37] produces false inner structures. The other approaches cannot represent open surfaces by nature.
@@ 5.1 Tessellation Speed
Remeshing [29] 1.208 0.427 0.941 0.178 0.025 1.498 0.264 0 0 11895 23785 25
DMTet [33] 0.923 0.251 0.926 0.228 0.017 6.466 0.001 0 0 18145 36315 201
FlexiCubes [34] 1.032 0.355 0.925 0.154 0.025 2.065 0.014 0 0.003 12570 24745 88
GShell [19] 0.542 0.402 0.939 0.150 0.077 2.784 0.063 0 0.002 11920 23137 208
DMesh [37] 0.494 0.393 0.937 0.116 0.060 1.799 0 0.032 0.001 2396 4863 751
DMesh++(Ours) 0.127 0.460 0.964 0.133 0.062 1.603 0 0 0 10021 19300 204
@@
Table 2: Quantitative comparison of 3D multi-view reconstruction results. We evaluate the reconstruction results with 5 metrics related to geometric accuracy and 4 metrics related to mesh quality. For geometric accuracy, we use Chamfer Distance (CD), F1 score (F1), Normal Consistency (NC), Edge Chamfer Distance (ECD), and Edge F1 score (EF1). For mesh quality, we use face Aspect Ratio (AR), Self-Intersection face Ratio (SI), Non-Manifold Edge ratio (NME), and Non-Manifold Vertex ratio (NMV). Finally, we additionally report several statistics: number of vertices and faces, and computation time. We highlighted the best results and the second best results for the evaluation metrics. For every criteria, DMesh++ achieves the best, or at least comparable results.
We compare the tessellation function speed between DMesh [37] (Eq. 1) and our DMesh++ (Eq. 2). For both 2D and 3D cases, we randomly generate $N$ points in a unit cube, find each point’s 10 nearest neighbors, and use these proximities to form potential face combinations. From these, we randomly select $N$ faces as query faces for the tessellation function. For each $N$ (ranging from $1K$ to $200K$ to reflect practical scenarios), we ran five trials and averaged the computational speeds.
In Fig. 6, we compare the computational costs of DMesh and DMesh++. For DMesh, costs increase linearly with point count in both 2D and 3D, due to the sequential WDT algorithm. In contrast, DMesh++ shows sub-linear scaling up to 50K points, benefiting from GPU parallelization (Sec. 3.2). Beyond this, costs rise more sharply due to GPU thread limitations. Nevertheless, DMesh++ handled 200K points in $117ms$ for 2D and $168ms$ for 3D. This demonstrates that the Minimum-Ball algorithm significantly accelerates tessellation, enabling efficient handling of complex shapes (Figs. 1 and 7) that DMesh cannot handle. See Appendix 13.1 for more results on complex 2D drawings.
5.2 Adaptive Resolution
We now demonstrate the impact of the Reinforce-Ball algorithm on 2D point cloud reconstruction for a font dataset. This experiment uses vector graphics of 26 uppercase letters from four different font styles. For each letter, we sampled a dense point cloud, which helps prevent “holes” that can occur with insufficient point density (Appendix 12.1).
In Tab. 1, we present quantitative ablation studies on the Reinforce-Ball algorithm. Increasing the tunable hyperparameter $\epsilon_{card}$ (Sec. 4.2), which controls regularization strength, leads to a rapid reduction in vertices and edges. For instance, with $\epsilon_{card}=10^{-5}$, edges decrease by nearly 94% with minimal impact on reconstruction quality. DMesh [37] also offers a tunable parameter, $\lambda_{weight}$, for weight regularization to reduce mesh complexity. However, while edge reduction occurs, DMesh’s reconstruction quality degrades more quickly. At $\epsilon_{card}=10^{-5}$, our method achieves a similar CD loss to DMesh with $\lambda_{weight}=10^{-4}$ but uses about 78% fewer edges. This advantage is also evident in Fig. 8, where our Reinforce-Ball algorithm removes redundant edges effectively and adapts the mesh to local geometry. In contrast, DMesh’s edge removal disregards local geometry, resulting in loss of detail. The main limitation of Reinforce-Ball is its computational cost; while memory-efficient, it converges slowly. Accelerating this algorithm would be a valuable direction for future work.
5.3 3D Multi-View Reconstruction
In this task, we reconstruct a mesh from multi-view images of a target object. We render $(512\times 512)$ diffuse and depth maps of the ground truth object from 64 viewpoints (Fig. 4). We selected 10 closed and 10 open surfaces from Thingi10K dataset [46] for this experiment. We chose objects with minimal self-occlusions, as we need dense observations of an object to fully reconstruct it (Appendix 12.2). For comparison, we tested five other mesh reconstruction algorithms: Remeshing [2], DMTet [33], FlexiCubes [34], GShell [19], and DMesh [37], each with settings optimized for best quality, including post-processing steps such as visibility tests to remove false internal structures. We also aimed to produce meshes of similar complexity. See Appendix 12.2 for details.
In Tab. 2 and Figs. 9 and 10, we present quantitative and qualitative comparisons of the reconstruction results. For quantitative analysis, we assess geometrical accuracy and mesh quality using five and four metrics, respectively, and also report statistics on mesh complexity and computation time. The values represent averages across the entire dataset; in Appendix 13.2, we separate results for closed and open surfaces.
@@ 5.2 Adaptive Resolution
In Tab. 2, we observe that Remeshing, DMTet, and FlexiCubes have high CD errors, largely because they cannot represent open surfaces (Fig. 9). This limitation also explains why Remeshing and FlexiCubes are faster than other methods. Although Remeshing achieved the best AR and avoided non-manifoldness, it generated numerous self-intersections, particularly on open surfaces (Fig. 10).
@@
GShell improves CD loss by representing open surfaces through sub-surface extraction from closed templates. However, it struggles with complex open surfaces, as estimating the closed templates is challenging at the first place (Fig. 9). It also suffers from self-intersections (Fig. 10) and suboptimal AR. Compared to that, DMesh produced self-intersection free mesh with better AR and CD, as it can represent open surfaces. Also, it produced much simpler mesh than the other methods. However, it sometimes produced false inner structure due to occlusion (Fig. 9), and was not free from non-manifoldness. Its largest drawback was in computational cost, limiting its utility for fine-grained reconstructions.
DMesh++ addresses this issue with the Minimum-Ball algorithm (Sec. 3.2) and nearest neighbor caching (Appendix 7.4). It achieves the best or comparable results across all metrics, with computational costs similar to other methods. Importantly, DMesh++ guarantees meshes free of self-intersections (Fig. 10) and non-manifoldness, offering geometrically accurate and high-quality reconstructions at efficient computational costs.
Colors.
For 3D multi-view reconstruction, we jointly optimize per-point colors to recover a textured object or small scene. For a face $F=\{p_{1},p_{2},p_{3}\}$, the color of a point $p$ on $F$ is determined by interpolating the colors of $p_{1},p_{2},$ and $p_{3}$ using barycentric coordinates. In our experiments, we used a similar setup as before, substituting diffuse maps with high-resolution colored renderings (Appendix 12.2). We reconstructed textured meshes from several objects and small scenes in Objaverse [8] (Figs. 1 and 11), demonstrating our method’s ability to handle textured meshes and potential for real-world image-based reconstruction. More in Appendix 13.3.
6 Discussion
Figure 11: 3D scene reconstructions from multi-view images. For the left scene, we render the front view at the top right corner.
We presented DMesh++, a probabilistic approach for efficiently managing mesh connectivity. While primarily demonstrated for reconstruction task on synthetic dataset, DMesh++ can be applied across a range of applications.
First, we can extend the current reconstruction pipeline to create large scenes from real-world images. Although other representations, like NeRF [23] and Gaussian Splatting [14], yield high-quality reconstruction results, they lack precise geometric information. DMesh++, however, directly produces meshes that can be readily used in downstream applications, such as physics simulations. We also envision DMesh++ being used to train generative models to understand mesh connectivity. Since DMesh++ accommodates diverse mesh topologies, generative models trained on it could produce meshes with varied topologies. For example, it could be applied to generate or infer complex structures in Biometrics like DNA (Fig. 1).
Limitations and Future Directions. DMesh++ has several limitations to address for broader utility. First, although DMesh++ offers a way to eliminate non-manifoldness in a combinatorial sense, this does not ensure a favorable mesh topology, such as watertight mesh. Removing faces to resolve non-manifoldness often creates small holes in the mesh surface, which are undesirable. Second, we only implement reconstruction of DMesh++ from input images or point clouds. Depending purely on inputs, issues like self-occlusions in multi-view reconstruction can lead to floating artifacts (Appendix 12.2). To handle this, future works can introduce topological constraints to DMesh++ and utilize data-driven priors to facilitate the reconstruction.
Acknowledgement This work was supported in part by Barry Mersky and Capital One E-Nnovate Endowed Professorships, National Science Foundation, U.S. Army Research Office, and Meta Graduate Fellowship.

### v2

@@ 4 Reconstruction Process
Figure 7: Comparison of tessellation cost. Our method computes face probabilities up to 16 times faster in 2D and 32 times faster in 3D than DMesh [42], while using up to 96% less GPU memory in 2D and 75% less in 3D.
@@ 5 Experiments
This section presents our experimental results. First, we evaluate how Minimum-Ball enhances the computational efficiency of the tessellation function in both 2D and 3D. Next, we demonstrate the results of a point cloud reconstruction task in 2D and 3D as a practical application. Finally, we showcase an application in 3D multi-view reconstruction. Together, these results illustrate that our method is well-suited for downstream tasks involving complex shapes. Our main algorithms are implemented in PyTorch [35] and CUDA [30]. All experiments were conducted on a system with an AMD EPYC 7R32 CPU and an NVIDIA A10 GPU. For details, refer to Appendix 9.
5.1 Tessellation Cost
We compare the computational cost of the tessellation function for DMesh [42] (Eq. 1) and our DMesh++ (Eq. 2). For both 2D and 3D scenarios, we randomly generate $N$ points within a unit cube, find each point’s 10 nearest neighbors, and use these proximities to form potential face combinations. From these, we randomly select $N$ faces as query faces for the tessellation function. For each value of $N$ (ranging from $1K$ to $200K$ to reflect practical scenarios), we conducted 5 trials and averaged the computational costs.
In Fig. 7, we compare the computational costs of DMesh and DMesh++. In terms of speed, DMesh’s performance scales linearly with the number of points in both 2D and 3D due to its sequential WDT algorithm. In contrast, DMesh++ exhibits sub-linear scaling up to 50K points, benefiting from GPU parallelization (Sec. 3.2). Beyond this range, computational costs increase more sharply because of GPU thread limitations; however, DMesh++ still processed 200K points in $117ms$ for 2D and $168ms$ for 3D. Regarding GPU memory usage, both methods scale linearly, but DMesh++ uses significantly less memory since it stores only the additional information related to the minimum balls, whereas DMesh must store all details of the power diagram on the GPU [^4] Note that the 2D version of DMesh requires more memory than the 3D version, as the 2D implementation is not as optimized using CUDA..
These results demonstrate that the Minimum-Ball algorithm significantly enhances tessellation efficiency, enabling the effective handling of complex shapes.
5.2 Reconstruction Tasks
Figure 8: Qualitative comparison of 2D point cloud reconstruction results. The outputs of DMesh [42] and DMesh++ are rendered in red and blue, respectively.
Method CD(×$\times 10^{-6}$$\downarrow$ # Verts. # Edges. Time (sec)
DMesh [42] 1.97 2506 2245 30.39
DMesh++ 1.82 2862 2793 11.33
Table 1: Quantitative comparison of the 2D point cloud reconstruction task for the font dataset. DMesh++ reconstructs 2D meshes with greater accuracy and efficiency than DMesh.
@@ 5.2 Reconstruction Tasks
VoroMesh [26] 19.591 0.352 0.855 0.054 0.072 145.7 0 0.001 0 64561 129338 11
DMesh++ 0.034 0.471 0.919 0.063 0.094 1.765 0 0.130 0.003 25415 58537 282
PSR [17] 10.164 0.392 0.943 0.302 0.026 5.218 0 0 0 139857 279739 4
PoNQ [27] 1.578 0.402 0.934 0.056 0.090 2.288 0 0.002 0 47254 94664 32
DMesh [42] 0.154 0.289 0.921 0.077 0.069 1.961 0 0.103 0.002 5815 13088 1147
DMesh++ 0.033 0.480 0.938 0.060 0.116 1.814 0 0.087 0.004 25396 55546 282
Table 2: Quantitative comparison of 3D point cloud reconstruction results over 50 manually chosen models from Thingi10K [52] and Objaverse [10]. We highlight the best results for each metric. The upper panel compares methods that use unoriented point clouds (VoroMesh, DMesh++), while the lower panel displays methods that use oriented point clouds (PSR, PoNQ, DMesh, DMesh++).
Dataset.
For the 2D point cloud reconstruction task, we used vector graphics of 26 letters from four different font styles downloaded from the Google Fonts service [^5] https://fonts.google.com/. Additionally, we employed six vector graphic images representing complex drawings from Adobe Stock [^6] https://stock.adobe.com/.
For the 3D reconstruction tasks, we randomly selected 500 models from Thingi10K [52] dataset, half of which are closed-surface models and the other half are open-surface models. Also, we manually chose 20 and 30 models from Thingi10K and Objaverse [10] dataset for better diversity. We normalized the model scale to $1$ before reconstruction.
Metrics.
@@
For 2D results, we report the Chamfer Distance (CD) of the reconstructed outputs for quantitative comparison. For 3D, we assess geometric accuracy using five metrics—CD (Chamfer Distance), F1 (F1 score), NC (Normal Consistency), ECD (Edge Chamfer Distance), and EF1 (Edge F1 score)—following [7]. Additionally, we evaluate mesh quality using four metrics: AR (Aspect Ratio), SI (Self-Intersection ratio), NME (Non-Manifold Edge ratio), and NMV (Non-Manifold Vertex ratio). Finally, we provide statistics on mesh complexity (e.g., vertex and face counts) along with the reconstruction time.
5.2.1 2D Point Cloud Reconstruction
Figure 9: Qualitative comparison of 3D point cloud reconstruction results for a toad sitting on a leaf. For each result, we render its diffuse image on the left, and view-point normals on the right.
In this task, we aim to reconstruct a 2D mesh from a 2D point cloud to demonstrate that the DMesh++ formulation is applicable in 2D. We primarily compare our reconstruction results on a font dataset with those obtained by DMesh [42], whose formulation is also easily extendable to 2D. As input, we sampled 1,000 points from each spline curve composing the font and downsampled the entire point cloud using a grid with a cell size of 0.005.
In Tab. 1, we present a quantitative comparison of the reconstruction results. We observe that DMesh++ reconstructs 2D meshes more faithfully than DMesh in terms of CD loss, while also running 2.6×2.6$\times$2.6 × faster. In Fig. 8, we render the reconstructed 2D meshes for two examples. The results show that DMesh++ produces significantly fewer holes compared to DMesh, which is consistent with the CD loss comparison. However, some of these holes are inevitable, as we mainly reconstruct our shape with CD loss and there are places that lack points.
Additionally, we reconstructed several complex 2D drawings from point clouds to demonstrate the computational efficiency of DMesh++. Since these drawings contain finer details than the fonts, we downsampled the entire point cloud using a grid with a cell size of $0.001$. When we attempted to reconstruct these drawings with DMesh, the GPU memory consumption became prohibitively high, resulting in an error during reconstruction. Therefore, for qualitative evaluation, we report only the results of DMesh++ in Fig. 1 and Appendix 9.2.2.
Before moving on, we’d like to introduce an experimental algorithm that is applicable to 2D mesh optimization called the Reinforce-Ball algorithm, which could be used for producing an efficient mesh that adapts to local geometry. Please see Appendix 10 for more details.
[…]
PSR [17] 12.1 / 16.3 0.47 / 0.45 0.97 / 0.96 0.45 / 0.37 0.01 / 0.01
PoNQ [27] 2.44 / 7.86 0.48 / 0.45 0.97 / 0.95 0.01 / 0.04 0.26 / 0.22
DMesh [42] 2.82 / 3.29 0.25 / 0.22 0.94 / 0.93 0.01 / 0.01 0.15 / 0.12
DMesh++ 0.37 / 0.37 0.48 / 0.47 0.96 / 0.95 0.02 / 0.02 0.25 / 0.23
Table 3: Quantitative comparison of 3D point cloud reconstruction results over 500 randomly chosen models from Thingi10K [52]. The results over closed surfaces and open surfaces are shown together: (closed / open). We highlight the best results for two different categories with red and blue, respectively.
5.2.2 3D Point Cloud Reconstruction
Figure 10: Qualitative comparison of 3D multi-view reconstruction results for open surface. Here we illustrate from back of an open surface model (the front view is rendered at the left top of (a)). Colors represent inside and outside facing surfaces.
Geometric Accuacy Mesh Quality Statistics
[…]
Remeshing [33] 0.977 0.343 0.907 0.088 0.044 1.562 0.220 0 0 13273 26540 25
DMTet [38] 1.395 0.191 0.868 0.145 0.032 6.175 0 0 0 20549 41131 201
FlexiCubes [39] 3.493 0.290 0.880 0.091 0.038 2.093 0.011 0 0.005 14811 28882 88
GShell [23] 5.807 0.326 0.904 0.116 0.047 2.793 0.047 0 0.004 14587 28054 209
DMesh [42] 0.697 0.328 0.898 0.104 0.045 1.820 0 0.042 0.002 2461 5058 772
DMesh++ 0.342 0.360 0.923 0.074 0.059 1.639 0 0.025 0.036 12507 23727 205
Table 4: Quantitative comparison of 3D multi-view reconstruction results over 50 manually chosen models from Thingi10K [52] and Objaverse [10]. We highlight the best results and the second best results for each evaluation metric.
In this task, we reconstruct a 3D mesh from a dense 3D point cloud. As input, we sampled 200K points from the ground truth mesh using the Poisson disk sampling algorithm [3] implemented in MeshLab [9]. For comparison, we employed other widely used optimization-based point cloud reconstruction methods, including Screened Poisson Surface Reconstruction (PSR)[17], VoroMesh[26], PoNQ [27], and DMesh [42]. For PSR, DMesh, and PoNQ, we used oriented point clouds for reconstruction, while for DMesh++ we tested both unoriented and oriented point clouds. We used the PSR implementation available in MeshLab. For DMesh, we used its default settings; for VoroMesh and PoNQ, we employed a grid size of 128 and optimized for 1000 epochs to achieve the best results.
We report the quantitative comparisons in Tabs. 2 and 3. In Tab. 2, results are averaged over the 50 manually chosen models and split into two categories: methods using unoriented versus oriented point clouds. In both categories, DMesh++ performs best or is comparable to other methods in geometric accuracy and triangle aspect ratio. Compared to DMesh, DMesh++ outperforms all metrics while handling 4.2$\times$ more faces with a 76% reduction in computation time. Furthermore, DMesh++ achieves significantly better CD, F1, and EF1 scores than PSR, VoroMesh, and PoNQ, which struggle with open surfaces. Separate evaluations in Tab. 3 over 500 randomly chosen models confirm that DMesh++ largely outperforms on open surfaces and performs comparably or better on closed surfaces.
These quantitative results align with the qualitative comparison in Fig. 9, where a toad sitting on a thin leaf is reconstructed . While PSR and VoroMesh fail to reconstruct the open surface of the leaf and PoNQ produces many holes, DMesh captures the overall geometry but misses fine details. DMesh++ successfully recovers both the overall shape and the intricate details. For qualitative comparison on a closed surface, see Fig. 19.
5.2.3 3D Multi-View Reconstruction
In this task, we reconstruct a mesh from multi-view images of a target object, assuming full knowledge of the rendering model and lighting conditions. Specifically, we employ simple Phong shading [36] with a directional light from the camera to the object for rendering the ground truth images. We generate $(512\times 512)$ diffuse and depth maps of the object from 64 viewpoints (Fig. 20) to supervise the reconstruction process. Note that we omit colors and textures in this experiment to focus solely on geometric quality.
[…]
REM [33] 58.0 / 27.5 0.33 / 0.32 0.90 / 0.90 0.06 / 0.07 0.11 / 0.09
DMT [38] 36.2 / 65.1 0.21 / 0.21 0.89 / 0.88 0.17 / 0.13 0.03 / 0.03
FLE [39] 14.5 / 34.1 0.37 / 0.35 0.92 / 0.91 0.04 / 0.06 0.10 / 0.08
GSH [23] 5.74 / 54.6 0.38 / 0.36 0.94 / 0.93 0.04 / 0.06 0.15 / 0.12
DME [42] 11.9 / 11.7 0.18 / 0.20 0.87 / 0.87 0.06 / 0.08 0.04 / 0.04
DMesh++ 3.35 / 3.82 0.37 / 0.36 0.94 / 0.93 0.03 / 0.03 0.21 / 0.17
Table 5: Quantitative comparison of 3D multi-view reconstruction results over 500 randomly chosen models from Thingi10K [52]. The results over closed surfaces and open surfaces are shown together: (closed / open). We highlight the best results for two different categories with red and blue, respectively.
For comparison, we evaluated five mesh reconstruction algorithms: Remeshing [2], DMTet [38], FlexiCubes [39], GShell [23], and DMesh [42]. We optimized each method to produce the best-quality meshes with similar vertex and face counts. Specifically, we set the grid sizes to 128 for DMTet, and 80 for both FlexiCubes and GShell. For more details, please refer to Appendix 9.4.
In Tabs. 4 and 5 and Figs. 10 and 11, we present quantitative and qualitative comparisons of the reconstruction results. In Tab. 4, we observe that Remeshing, DMTet, and FlexiCubes have high CD errors, largely because they cannot represent open surfaces (Fig. 10). This limitation also explains why Remeshing and FlexiCubes are faster than other methods. Although Remeshing achieved the best AR and avoided non-manifoldness, it generated numerous self-intersections, particularly on open surfaces (Fig. 11).
@@
When it comes to GShell, we found out that it usually improves CD loss by representing open surfaces through sub-surface extraction from closed templates. However, it still struggles with complex open surfaces (Fig. 10). Also, it had a robustness issue with several outlier models – which is the main reason of high average CD, even though it achieved much better CD for most of the models. Additionally, it also suffers from self-intersections (Fig. 11) and suboptimal AR. Compared to that, DMesh produced self-intersection free mesh with better CD and AR, as it can represent open surfaces robustly. Also, it produced much simpler mesh than the other methods. However, it sometimes produced false inner structure due to occlusion (Fig. 10), and its largest drawback was in computational cost, limiting its utility for fine-grained reconstructions.
DMesh++ addresses this issue using the Minimum-Ball algorithm (Sec. 3.2) and nearest neighbor caching (Appendix 7.4). It achieves the best or comparable results across all metrics while maintaining computational costs similar to those of other methods. Furthermore, as shown in Tab. 5, DMesh++ delivers superior, or at least comparable performance for both closed and open surfaces. Qualitative comparisons in Figs. 10 and 20 also support this observation. These results prove the robustness of DMesh++ in handling complex shapes with diverse topologies, whereas other methods are limited in one aspect or another.
Colors. As mentioned in Fig. 2 and Sec. 3.1, points can carry additional features. Here, we demonstrate that we can jointly optimize per-point colors to recover a textured shape from multi-view images. For a point on a face, the color of the point is determined by barycentric interpolation of the face vertex colors. Under the assumption that we know all the rendering and lighting conditions, we can reconstruct textured meshes from multi-view images as shown in Figs. 1, 6 and 12. In particular, Fig. 12 demonstrates that DMesh++ can reconstruct a small scene on which physics simulations, such as bouncing balls, can be run directly. These results highlight a promising future direction of jointly optimizing additional per-point features.
6 Conclusion
We presented DMesh++, a probabilistic approach for efficient, differentiable mesh connectivity handling. Our Minimum-Ball algorithm significantly reduces computational costs, enabling DMesh++ to recover 2D and 3D shapes with diverse topologies from point clouds or multi-view images more effectively than baseline methods.
Limitations. There are areas where DMesh++ can be further improved. First, some non-manifoldness errors remain in the reconstruction results (Tabs. 2 and 4), while observing significant improvement over DMesh [42]. We conjecture that there is a trade-off between the representation power and the occurrence of these topological errors; more careful analysis of the trade-off is needed to maintain expressiveness while eliminating errors. Second, there are application-specific challenges. Although our method yields superior 3D point cloud reconstruction results, it incurs higher computational costs (Tab. 2), and it is less effective for sparse point clouds. In 3D multi-view reconstruction, we cannot use the current implementation for real-world images, as discussed in Appendix 9.4.2.
Figure 12: Physics simulation on a staircase reconstructed from multi-view images. We simulate the motion of bouncing balls directly on the mesh generated by DMesh++.
Future Directions. For 3D point cloud reconstruction, we plan to explore alternative loss formulations to CD loss, as its computation currently dominates the overall cost.
To extend our 3D multi-view reconstruction algorithm to real-world images, we will integrate DMesh++ with alternative representations such as Gaussian Splatting (GS) [18, 13]. By appending GS features to the per-point features and optimizing them jointly, we expect to recover high-quality meshes that we can readily use for downstream tasks.
Furthermore, we envision leveraging DMesh++ to train generative models that capture complex mesh connectivity. Currently, the Minimum-Ball algorithm only uses point positions for connectivity. In future work, we plan to incorporate additional per-point features into this algorithm to train generative models capable of understanding diverse and intricate mesh connectivity, such as that of DNA (Fig. 1).
We believe our work establishes an important foundation for harnessing differentiable, probabilistic mesh within the current optimization framework, and we hope it will further drive future downstream applications.